# Explain Like I’m Five: Using LLMs to Improve PDE Surrogate Models with Text

### v2

@@ 5.2 Autoregressive Rollout
Second, we perform autoregressive rollout starting from our initial condition for 40 steps. This evaluates how well our model is able to temporally extrapolate beyond our training horizon. In this case, we use the models trained in section 5.1 with no additional fine-tuning. Accumulated rollout error is presented in table 1. We see that our SentenceTransformer FactFormer has lowest accumulated error for four of five data sets both for standard training and transfer learning, and outperforms baseline across all data sets. Our Llama FactFormer also outperforms baseline across all of our data sets except Shallow Water with standard training. For SentenceTransformer Factformer, we have an average reduction in accumulated error of 64.2% and 84.7% across all of our data set for standard training and transfer learning, respectively. For Llama FactFormer, we have an average reduction in accumulated error of 6.2% and 81.2% for standard and transfer learning, respectively. When removing Shallow Water for Llama Factformer with standard training, we have an average reduction in error of 58.0%.
@@
Combined 1.36 $\pm$ 0.03 0.78 $\pm$ 0.04 0.90 $\pm$ 0.02 3.89 $\pm$ 0.01 0.80 $\pm$ 0.09 0.86 $\pm$ 0.04
Burgers 3.49 $\pm$ 0.15 0.97 $\pm$ 0.10 1.41 $\pm$ 0.42 12.17 $\pm$ 0.13 0.85 $\pm$ 0.07 1.03 $\pm$ 0.10
Heat 1.08 $\pm$ 0.10 0.16 $\pm$ 0.01 0.47 $\pm$ 0.09 4.61 $\pm$ 0.03 0.30 $\pm$ 0.17 0.13 $\pm$ 0.07
Navier Stokes 1.52 $\pm$ 0.46 0.54 $\pm$ 0.25 0.26 $\pm$ 0.09 2.56 $\pm$ 0.03 0.25 $\pm$ 0.00 0.40 $\pm$ 0.08
Shallow Water 0.12 $\pm$ 0.09 0.05 $\pm$ 0.04 0.35 $\pm$ 0.36 156.60 $\pm$ 1.64 51.22 $\pm$ 60.16 65.70 $\pm$ 21.50
@@ 5.4 Ablation Study
Combined 3.89 $\pm$ 0.01 0.84 $\pm$ 0.02 0.85 $\pm$ 0.01 0.85 $\pm$ 0.08 0.87 $\pm$ 0.01 0.82 $\pm$ 0.07 0.80 $\pm$ 0.02 0.82 $\pm$ 0.07 0.80 $\pm$ 0.09
Burgers 12.17 $\pm$ 0.13 1.02 $\pm$ 0.04 0.86 $\pm$ 0.08 1.10 $\pm$ 0.05 1.04 $\pm$ 0.07 0.86 $\pm$ 0.04 0.86 $\pm$ 0.09 1.00 $\pm$ 0.07 0.85 $\pm$ 0.07
Heat 4.61 $\pm$ 0.03 0.37 $\pm$ 0.02 0.30 $\pm$ 0.19 0.38 $\pm$ 0.03 0.38 $\pm$ 0.05 0.20 $\pm$ 0.04 0.19 $\pm$ 0.03 0.38 $\pm$ 0.02 0.30 $\pm$ 0.17
Navier Stokes 2.56 $\pm$ 0.03 0.19 $\pm$ 0.15 0.29 $\pm$ 0.04 0.26 $\pm$ 0.02 0.33 $\pm$ 0.06 0.27 $\pm$ 0.02 0.27 $\pm$ 0.01 0.24 $\pm$ 0.03 0.25 $\pm$ 0.00
Shallow Water 156.60 $\pm$ 1.64 82.06 $\pm$ 33.18 10.02 $\pm$ 9.98 56.15 $\pm$ 50.17 51.89 $\pm$ 33.97 63.49 $\pm$ 50.75 43.81 $\pm$ 50.04 65.07 $\pm$ 75.05 51.22 $\pm$ 60.16

### v3

@@ 5.2 Autoregressive Rollout
Second, we perform autoregressive rollout starting from our initial condition for 40 steps. This evaluates how well our model is able to temporally extrapolate beyond our training horizon. In this case, we use the models trained in section 5.1 with no additional fine-tuning. Accumulated rollout error is presented in table 1. We see that our SentenceTransformer FactFormer has lowest accumulated error for four of five data sets both for standard training and transfer learning, and outperforms baseline across all data sets. Our Llama FactFormer also outperforms baseline across all of our data sets except Shallow Water with standard training. For SentenceTransformer Factformer, we have an average reduction in accumulated error of 64.2% and 54.2% across all of our data set for standard training and transfer learning, respectively. For Llama FactFormer, we have an average reduction in accumulated error of 6.2% and 45.3% for standard and transfer learning, respectively. When removing Shallow Water for Llama Factformer with standard training, we have an average reduction in error of 63.1%.
@@
Combined 1.36 $\pm$ 0.03 0.78 $\pm$ 0.04 0.90 $\pm$ 0.02 1.30 $\pm$ 0.08 0.80 $\pm$ 0.09 0.86 $\pm$ 0.04
Burgers 3.49 $\pm$ 0.15 0.97 $\pm$ 0.10 1.41 $\pm$ 0.42 4.06 $\pm$ 3.93 0.85 $\pm$ 0.07 1.03 $\pm$ 0.10
Heat 1.08 $\pm$ 0.10 0.16 $\pm$ 0.01 0.47 $\pm$ 0.09 1.54 $\pm$ 0.27 0.30 $\pm$ 0.17 0.13 $\pm$ 0.07
Navier Stokes 1.52 $\pm$ 0.46 0.54 $\pm$ 0.25 0.26 $\pm$ 0.09 0.85 $\pm$ 0.57 0.25 $\pm$ 0.00 0.40 $\pm$ 0.08
Shallow Water 0.12 $\pm$ 0.09 0.05 $\pm$ 0.04 0.35 $\pm$ 0.36 52.20 $\pm$ 50.63 51.22 $\pm$ 60.16 65.70 $\pm$ 21.50
@@ 5.4 Ablation Study
Combined 1.30 $\pm$ 0.08 0.84 $\pm$ 0.02 0.85 $\pm$ 0.01 0.85 $\pm$ 0.08 0.87 $\pm$ 0.01 0.82 $\pm$ 0.07 0.80 $\pm$ 0.02 0.82 $\pm$ 0.07 0.80 $\pm$ 0.09
Burgers 4.06 $\pm$ 3.93 1.02 $\pm$ 0.04 0.86 $\pm$ 0.08 1.10 $\pm$ 0.05 1.04 $\pm$ 0.07 0.86 $\pm$ 0.04 0.86 $\pm$ 0.09 1.00 $\pm$ 0.07 0.85 $\pm$ 0.07
Heat 1.54 $\pm$ 0.27 0.37 $\pm$ 0.02 0.30 $\pm$ 0.19 0.38 $\pm$ 0.03 0.38 $\pm$ 0.05 0.20 $\pm$ 0.04 0.19 $\pm$ 0.03 0.38 $\pm$ 0.02 0.30 $\pm$ 0.17
Navier Stokes 0.85 $\pm$ 0.57 0.19 $\pm$ 0.15 0.29 $\pm$ 0.04 0.26 $\pm$ 0.02 0.33 $\pm$ 0.06 0.27 $\pm$ 0.02 0.27 $\pm$ 0.01 0.24 $\pm$ 0.03 0.25 $\pm$ 0.00
Shallow Water 52.20 $\pm$ 50.63 82.06 $\pm$ 33.18 10.02 $\pm$ 9.98 56.15 $\pm$ 50.17 51.89 $\pm$ 33.97 63.49 $\pm$ 50.75 43.81 $\pm$ 50.04 65.07 $\pm$ 75.05 51.22 $\pm$ 60.16